length of the rectangle is 2 inches more then the width. The area is 35 square inches. Find the dimensions
Answer by oberobic(2304)

L = W+2
A = L*W = 35
.
substitute
.
(W+2)*W = 35
W^2 +2W -35 = 0
.
factor
.
(W+7)(W-5) = 0
.
W = -7 or 5
.
Negative width does not make sense, so we select W =5.
L = W+2= 7
.
Answer: The rectangle has width = 5 inches and length = 7 inches.
.
Done.